# Scalable Decentralized Algorithms for Online Personalized Mean Estimation

### v3

@@ D.4 $(\varepsilon,\delta)$-Bounds: Proof of Theorem 7
$\prescript{}{c}{\tau}^{C}=\max\left\{\zeta_{D},C^{\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\mathbf{x}-\bm{\mu}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\left(\frac{(1-\alpha)^{2}}{(1-\alpha\lambda_{2}(W))^{2}}\left(1+\frac{1}{\ln{\frac{1}{\alpha\lambda_{2}(W)}}}\right)^{2}+\frac{1}{n_{c}^{2}}\left(1+\frac{1}{\ln{\frac{1}{\alpha}}}\right)^{2}\right)\right\}$
@@
$\prescript{}{c}{\tau}^{C}\triangleq\max\left\{\zeta_{D},C^{\prime\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\ln^{2}\left(eC^{\prime\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\right)\right\}.$
@@
We start considering the auxiliary system studied in the previous sections: consensus matrices can be arbitrary until time $\zeta_{D}$ and then agents acquire perfect knowledge about which neighbours belong to the same class and simply rely on information arriving through these links.
@@
$t^{\prime}\geqslant\prescript{}{c}{\tau}^{C}=\max\left\{\zeta_{D},C^{\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\mathbf{x}-\bm{\mu}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\left(\frac{(1-\alpha)^{2}}{(1-\alpha\lambda_{2}(W))^{2}}\left(1+\frac{1}{\ln{\frac{1}{\alpha\lambda_{2}(W)}}}\right)^{2}+\frac{1}{n_{c}^{2}}\left(1+\frac{1}{\ln{\frac{1}{\alpha}}}\right)^{2}\right)\right\}.$ (82)
@@
$\prescript{}{c}{\tau}^{C}\triangleq\max\left\{\zeta_{D},C^{\prime\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\ln^{2}\left(eC^{\prime\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\right)\right\}.$ (94)
Finally, let us consider the system of interest. With probability $21-\delta/$ 2 the agents will have identified the correct links by time $\zeta_{D}$. The corresponding trajectories coincide with trajectories of the auxiliary system we studied. For the auxiliary system, the estimates have the required precision after time $\prescript{}{c}{\tau}^{C}$ with probability $21-\delta/$ 2. It follows that the probability that the estimates in the stochastic have not the required precision after time $\prescript{}{c}{\tau}^{C}$ is at most $\delta$. ∎
Appendix E Appendix - On the Structure of $G(N,r)$ and its Impact on Performance of our Algorithms
E.1 On the local tree structure of $G(N,r)$
Let $\mathcal{N}=G(\mathcal{V},\mathcal{E})$ be a network sampled from the class of random regular graphs $G(N,r)$ with fixed degree $r$. 999observe that graphs in $G(N,r)$ are not necessarily simple Here and in the following we fix $\mathcal{V}=\mathcal{A}$ and $n=|\mathcal{V}|=|\mathcal{A}|$. We are interested to investigate the structure of the $d$-deep neighborhood, $\mathcal{N}^{(v)}_{d}$, of a generic node $v$ (i.e. the sub-graph inter-connecting nodes at distance smaller or equal than $d$v from $v$). Observe that as $n$ grows large, for any $d\in\mathbb{N}$, we should expect that $\mathcal{N}^{(v)}_{d}$ is likely equal to $\mathcal{T}_{d}$, a perfectly balanced tree of depth $d$, in which the root has $r$ children, and all the other nodes have $r-1$. children.
To this end using an approach inspired by Lemma 5 in Rossi et al. [2019] (which applies to directed graphs), we can claim that:
@@ E.1 On the local tree structure of $G(N,r)$
Proof First observe that realizations of $G(N,r)$ graphs are typically obtained through the following standard procedure: every node is initially connected to $r\geq 2$ stubs; then stubs then are sequentially randomly matched/paired to form edges, as follows: at each stage, an arbitrarily selected free/unpaired stub is selected and paired with a different stub picked uniformly at random among the still unpaired stubs.
@@
Our procedure explores the $d$-neighborhood of a node $v_{0}$, taken at random, by sequentially unveiling stub-pairings, according to a bread-first approach. At every step the procedure checks weather the already explored-portion of $\mathcal{N}^{(v_{0})}_{d}$, $G(\mathcal{V},\mathcal{E})$, has a tree structure.
@@
In case $v_{1}\neq v_{0}$ the algorithm proceeds, otherwise it prematurely terminates providing $G(\mathcal{V}^{(1)},\mathcal{E}^{(1)})$. At step 2, our procedure takes a new unmatched stub $k_{2}$ (connected again to $v_{0}$). $k_{2}$ is matched uniformly at random with another free (i.e. still unmatched) stub $r(k_{2})\not\in\{k_{1},r(k_{1}),k_{2}\}$, let $v_{2}:=\nu(r(k_{2}))$. Then sets $\mathcal{V}$ and $\mathcal{E}$ are updated: $\mathcal{V}^{(2)}=\mathcal{V}^{(1)}\cup\{v_{2}\}$ and $\mathcal{E}^{(2)}=\mathcal{E}^{(2)}\cup\{(v_{0},v_{2})\}$. $G(\mathcal{V}^{(2)},\mathcal{E}^{(2)})$ is a tree only under the condition that $v_{2}:=\nu(r(k_{2}))\not\in\mathcal{V}^{(1)}$. This happens with a probability:
@@
$M=\sum_{v\in\mathcal{V}}\mathbb{1}_{\{\mathcal{N}^{(v)}_{d}\neq\mathcal{T}_{d}\}}$
@@
$\mathbb{E}[M]$ $=\sum_{v\in\mathcal{V}}\mathbb{E}[\mathbb{1}_{\{\mathcal{N}{(v)}_{d}\neq\mathcal{T}_{d}\}}]=\sum_{v\in\mathcal{V}}\mathbb{P}(\mathcal{N}^{(v)}_{d}\neq\mathcal{T}_{d})$
@@
At last we would like to highlight that (95) can be transferred to the class $G_{0}(N,r)$ of uniformly chosen simple regular graphs thanks to:
Proposition 24 (Amini [2010]).
@@
At last we would like to mention the following result, from which Proposition 24 rather immediately descends.
Theorem 25 (Janson [2009]).
@@
Observe Theorem 25 provides a theoretical foundation to the to design a simple algorithm for the generation of a graph in class $G_{0}((N,r)$, based on the superposition of an acceptance/rejection procedure to the generation of graphs in $G(N,r)$. More efficient algorithms are, however, well known in literature McKay and Wormald [1990].
E.2 The structure of $\mathcal{CC}_{a}^{d}$ and $\mathcal{CC}_{a}$
@@
Proof
The proof is rather immediate. Consider $a$ and explore $\mathcal{CC}_{a}^{d}$ according to a breath-first exploration process that stops at depth-$d$. The number of off-springs of $a$, $O_{a}$, by construction, equals the number of nodes in $\mathcal{N}_{a}$ that belongs to class $k_{a}$. This number $O_{a}$ is distributed as: $O_{a}\stackrel{{\scriptstyle L}}{{=}}\text{Bin}(r,p_{k_{a}})$. Consider, now, any other explored node $a^{\prime}$, at distance $i<d$ from the ancestor, this node, by construction, will have a unique parent node $p_{a^{\prime}}$, Off-springs of $a^{\prime}$, $O_{a^{\prime}}$ will be given by all nodes in $\mathcal{N}_{a^{\prime}}\setminus\{p_{a^{\prime}}\}$ that belong to class $k_{a}$. Their number, $O_{a^{\prime}}$, is distributed as: $O_{a^{\prime}}\stackrel{{\scriptstyle L}}{{=}}\text{Bin}(r,p_{k_{a^{\prime}}}).$ See fig. 6. When the exploration reaches a $d$-depth node, it stops, therefore the number of off-springs for every $d$-depth nodes is zero. A last, since the sets $\mathcal{N}_{a^{\prime}}\setminus\{p_{a^{\prime}}\}$ are disjoint (as immediate consequence of the fact that $\mathcal{N}^{d}_{a}$ is a tree), the variables in $\{O_{a^{\prime}}\}_{a^{\prime}\in\mathcal{CC}_{a}^{d}}$ are independent.
@@
We can now recall and adapt a few standard results asymptotic results on Galton-Watson (GW) Processes (in a standard GW process all nodes in the tree give origin to number of off-springs, which are identically distributed and independent). First observe that, in our case, as $|\mathcal{A}|\to\infty$ we can assume that $d\to\infty$ as well, for example choosing $d$ as in Proposition 8.
We say that a standard GW process is super-critical if the average number off-springs of every node $\mathbb{E}[O_{a}]:=m>1$ Now for a standard supercritical GW process, denoted with $Z_{i}$ the number of nodes belonging to generation $i$, we have:
@@
For every super-critical nontrivial 101010 a GW process is non trivial if $\mathbb{P}(O_{a}=0)>0$ GW process $\{Z_{i}\}_{i}$ is bound to either extinction or explosion, i.e.,
@@
Previous result cam immediately extended to our two stages branching process as $d\to\infty$. Denoted with $m_{0}=rp_{k_{a}}$ and $m:=(r-1)p_{k_{a}}$, under the assumption that $m>1$ we obtain that:
@@
Indeed the distribution of off-springs at the root, has not impact on structural properties of the process.
@@
Consider a two-stages branching process, in which the number of offspring of the root is $\text{Bin}(p_{k},r)$ while the off-spring of every other node $\text{Bin}(p_{k},r-1)$. Its extinction probability $q_{2BP}$ is given by
$q_{2BP}=\sum_{h=0}^{r}{r\choose h}p_{k}^{h}(1-p_{k})^{r-h}q_{GW}^{h}=[(1-p_{k})+p_{k}q_{GW}]^{r}$
@@
. Then conditioning on the number of off-springs of the ancestor we get the claim. Of course previously computed asymptotic extinction probability $q_{2BP}$ provides an upper-bound to the probability of early extinction of a $d$-depth truncated two-stages Branching process.
@@
Appendix F Appendix - Proof of the results in Table 1

### v4

@@ F.4 $(\varepsilon,\delta)$-Bounds: Proof of Theorem 3
$\tau_{a}^{C}=\max\left\{\zeta_{D},C^{\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\mathbf{x}-\bm{\mu}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\left(\frac{(1-\alpha)^{2}}{(1-\alpha\lambda_{2}(W))^{2}}\left(1+\frac{1}{\ln{\frac{1}{\alpha\lambda_{2}(W)}}}\right)^{2}+\frac{1}{n_{c}^{2}}\left(1+\frac{1}{\ln{\frac{1}{\alpha}}}\right)^{2}\right)\right\}$
@@
$\tau_{a}^{C}\triangleq\max\left\{\zeta_{D},C^{\prime\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\ln^{2}\left(eC^{\prime\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\right)\right\}.$
@@
We start considering the auxiliary system studied in the previous sections: consensus matrices can be arbitrary until time $\zeta_{D}$ and then agents acquire perfect knowledge about which neighbors belong to the same class and simply rely on information arriving through these links.
@@
$t^{\prime}\geqslant\tau_{a}^{C}=\max\left\{\zeta_{D},C^{\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\mathbf{x}-\bm{\mu}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\left(\frac{(1-\alpha)^{2}}{(1-\alpha\lambda_{2}(W))^{2}}\left(1+\frac{1}{\ln{\frac{1}{\alpha\lambda_{2}(W)}}}\right)^{2}+\frac{1}{n_{c}^{2}}\left(1+\frac{1}{\ln{\frac{1}{\alpha}}}\right)^{2}\right)\right\}.$ (82)
@@
$\tau_{a}^{C}\triangleq\max\left\{\zeta_{D},C^{\prime\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\ln^{2}\left(eC^{\prime\prime}\frac{\operatorname*{\mathbb{E}}\left[\lVert\prescript{}{c}{P}\prescript{}{c}{\mathbf{x}}-\prescript{}{c}{\bm{\mu}}\rVert^{4}\right]}{n_{c}\varepsilon^{4}\delta}\right)\right\}.$ (94)
Finally, let us consider the system of interest. With probability $21-\delta/$ 2 the agents will have identified the correct links by time $\zeta_{D}$. The corresponding trajectories coincide with trajectories of the auxiliary system we studied. For the auxiliary system, the estimates have the required precision after time $\tau_{a}^{C}$ with probability $21-\delta/$ 2. It follows that the probability that the estimates in the stochastic have not the required precision after time $\tau_{a}^{C}$ is at most $\delta$. ∎
Appendix G Appendix G - On the Structure of $G(N,r)$ and its Impact on Performance of our Algorithms
G.1 On the Local Tree Structure of $G(N,r)$
Let $\mathcal{N}=G(\mathcal{V},\mathcal{E})$ be a network sampled from the class of random regular graphs $G(N,r)$ with fixed degree $r$. 666observe that graphs in $G(N,r)$ are not necessarily simple Here and in the following we fix $\mathcal{V}=\mathcal{A}$ and $n=|\mathcal{V}|=|\mathcal{A}|$. We are interested to investigate the structure of the $d$-deep neighborhood, $\mathcal{N}^{(v)}_{d}$, of a generic node $v$ (i.e. the sub-graph inter-connecting nodes at a distance smaller or equal than $d$v from $v$). Observe that as $n$ grows large, for any $d\in\mathbb{N}$, we should expect that $\mathcal{N}^{(v)}_{d}$ is likely equal to $\mathcal{T}_{d}$, a perfectly balanced tree of depth $d$, in which the root has $r$ children, and all the other nodes have $r-1$. children.
To this end using an approach inspired by Lemma 5 in (Rossi, Como, and Fagnani 2019) (which applies to directed graphs), we can claim that:
@@ G.1 On the Local Tree Structure of $G(N,r)$
Proof.
First observe that realizations of $G(N,r)$ graphs are typically obtained through the following standard procedure: every node is initially connected to $r\geq 2$ stubs; then stubs then are sequentially randomly matched/paired to form edges, as follows: at each stage, an arbitrarily selected free/unpaired stub is selected and paired with a different stub picked uniformly at random among the still unpaired stubs.
@@
Our procedure explores the $d$-neighborhood of a node $v_{0}$, taken at random, by sequentially unveiling stub-pairings, according to a bread-first approach. At every step, the procedure checks whether the already explored portion of $\mathcal{N}^{(v_{0})}_{d}$, $G(\mathcal{V},\mathcal{E})$, has a tree structure.
@@
In case $v_{1}\neq v_{0}$ the algorithm proceeds, otherwise it prematurely terminates providing $G(\mathcal{V}^{(1)},\mathcal{E}^{(1)})$. In step 2, our procedure takes a new unmatched stub $k_{2}$ (connected again to $v_{0}$). $k_{2}$ is matched uniformly at random with another free (i.e. still unmatched) stub $r(k_{2})\not\in\{k_{1},r(k_{1}),k_{2}\}$, let $v_{2}:=\nu(r(k_{2}))$. Then sets $\mathcal{V}$ and $\mathcal{E}$ are updated: $\mathcal{V}^{(2)}=\mathcal{V}^{(1)}\cup\{v_{2}\}$ and $\mathcal{E}^{(2)}=\mathcal{E}^{(2)}\cup\{(v_{0},v_{2})\}$. $G(\mathcal{V}^{(2)},\mathcal{E}^{(2)})$ is a tree only under the condition that $v_{2}:=\nu(r(k_{2}))\not\in\mathcal{V}^{(1)}$. This happens with a probability:
@@
$M=\sum_{v\in\mathcal{V}}\bm{1}_{\{\mathcal{N}^{(v)}_{d}\neq\mathcal{T}_{d}\}}$
@@
$\mathbb{E}[M]$ $=\sum_{v\in\mathcal{V}}\mathbb{E}[\bm{1}_{\{\mathcal{N}{(v)}_{d}\neq\mathcal{T}_{d}\}}]=\sum_{v\in\mathcal{V}}\mathbb{P}(\mathcal{N}^{(v)}_{d}\neq\mathcal{T}_{d})$
@@
At last, we would like to highlight that (95) can be transferred to the class $G_{0}(N,r)$ of uniformly chosen simple regular graphs thanks to:
Proposition 24 (Amini (2010)).
@@
At last, we would like to mention the following result, from which Proposition 24 rather immediately descends.
Theorem 25 ((Janson 2009)).
@@
Observe Theorem 25 provides a theoretical foundation to design a simple algorithm for the generation of a graph in class $G_{0}((N,r)$, based on the superposition of an acceptance/rejection procedure to the generation of graphs in $G(N,r)$. More efficient algorithms are, however, well-known in literature (McKay and Wormald 1990).
G.2 The Structure of $\mathcal{CC}_{a}^{d}$ and $\mathcal{CC}_{a}$
@@
Proof.
The proof is rather immediate. Consider $a$ and explore $\mathcal{CC}_{a}^{d}$ according to a breath-first exploration process that stops at depth-$d$. The number of off-springs of $a$, $O_{a}$, by construction, equals the number of nodes in $\mathcal{N}_{a}$ that belongs to class $k_{a}$. This number $O_{a}$ is distributed as: $O_{a}\stackrel{{\scriptstyle L}}{{=}}\text{Bin}(r,p_{k_{a}})$. Consider, now, any other explored node $a^{\prime}$, at distance $i<d$ from the ancestor, this node, by construction, will have a unique parent node $p_{a^{\prime}}$, off-springs of $a^{\prime}$, $O_{a^{\prime}}$ will be given by all nodes in $\mathcal{N}_{a^{\prime}}\setminus\{p_{a^{\prime}}\}$ that belong to class $k_{a}$. Their number, $O_{a^{\prime}}$, is distributed as: $O_{a^{\prime}}\stackrel{{\scriptstyle L}}{{=}}\text{Bin}(r,p_{k_{a^{\prime}}}).$ See Fig. 7. When the exploration reaches a $d$-depth node, it stops, therefore the number of off-springs for every $d$-depth nodes is zero. A last, since the sets $\mathcal{N}_{a^{\prime}}\setminus\{p_{a^{\prime}}\}$ are disjoint (as immediate consequence of the fact that $\mathcal{N}^{d}_{a}$ is a tree), the variables in $\{O_{a^{\prime}}\}_{a^{\prime}\in\mathcal{CC}_{a}^{d}}$ are independent. ∎
@@
We can now recall and adapt a few standard asymptotic results on Galton-Watson (GW) Processes (in a standard GW process all nodes in the tree give origin to number of off-springs, which are identically distributed and independent). First observe that, in our case, as $|\mathcal{A}|\to\infty$ we can assume that $d\to\infty$ as well, for example choosing $d$ as in Proposition 6.
We say that a standard GW process is super-critical if the average number of off-springs of every node $\mathbb{E}[O_{a}]:=m>1$. Now for a standard supercritical GW process, denoted with $Z_{i}$ the number of nodes belonging to generation $i$, we have:
@@
For every super-critical nontrivial 777 a GW process is non trivial if $\mathbb{P}(O_{a}=0)>0$ GW process $\{Z_{i}\}_{i}$ is bound to either extinction or explosion, i.e.,
@@
The previous result can immediately extended to our two-stage branching process as $d\to\infty$. Denoted with $m_{0}=rp_{k_{a}}$ and $m:=(r-1)p_{k_{a}}$, under the assumption that $m>1$ we obtain that:
@@
Indeed the distribution of off-springs at the root has no impact on the structural properties of the process.
@@
Consider a two-stage branching process, in which the number of offspring of the root is $\text{Bin}(p_{k},r)$ while the off-spring of every other node $\text{Bin}(p_{k},r-1)$. Its extinction probability $q_{2BP}$ is given by
$q_{2BP}=\sum_{h=0}^{r}\binom{r}{h}p_{k}^{h}(1-p_{k})^{r-h}q_{GW}^{h}=[(1-p_{k})+p_{k}q_{GW}]^{r}$
@@
. Then conditioning on the number of off-springs of the ancestor we get the claim. Of course, previously computed asymptotic extinction probability $q_{2BP}$ provides an upper bound to the probability of early extinction of a $d$-depth truncated two-stage Branching process. ∎
@@
Appendix H Appendix H - Proof of the results in Table 2